# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ I-A Background
The increasing prevalence of Internet of things (IoT) devices and the demand for real-time applications have propelled a shift from centralized cloud computing to a distributed edge-cloud paradigm [1, 2, 3]. While cloud computing initially offers scalability and cost-effectiveness, limitations such as latency and bandwidth constraints spur the emergence of edge computing, bringing computation closer to data sources [4, 5, 6]. Consequently, edge-cloud collaborative computing (ECCC) arises, integrating the strengths of both paradigms to facilitate efficient resource utilization and enable complex AI tasks across the edge-cloud continuum [7, 8]. AI integration within edge-cloud systems addresses the inherent complexity of distributed environments, enabling optimized resource allocation and strategic model deployment for enhanced system performance [9]. In addition, AI empowers intelligent decision-making at the edge, enabling real-time responses. However, realizing distributed intelligence in edge-cloud systems introduces challenges, including efficient model partitioning and resource-aware model optimization, along with ensuring seamless edge-cloud collaboration and addressing security and privacy concerns [10, 11].
@@
Edge-cloud collaborative computing has emerged as a crucial paradigm to support the growing demands of real-time, intelligent applications, particularly with the proliferation of artificial intelligence of things (AIoT) devices [12]. However, inherent limitations in edge devices, such as constrained computational power and limited battery life, hinder their ability to independently handle computationally intensive AI workloads [13]. Additionally, the dynamic nature of workloads in these environments requires sophisticated resource allocation and task offloading strategies to maintain performance and prevent bottlenecks [14]. Applications like autonomous driving [15] and industrial automation [16] further necessitate low latency and high reliability, demanding intelligent resource management and optimization [17]. Consequently, AI offers a powerful set of tools, including deep reinforcement learning (DRL) for dynamic resource allocation [18] and multi-objective optimization for task scheduling [19], to address these challenges and improve the efficiency and performance of edge-cloud systems. For instance, federated learning enables collaborative model training across distributed devices while preserving privacy [13], ultimately contributing to enhanced quality of service (QoS) by optimizing resource utilization and minimizing latency.
@@
The evolution of communication networks plays a pivotal role in enabling effective edge-cloud collaboration by providing the foundational infrastructure that connects distributed computational resources. Recent advancements in network technologies, particularly software-defined networking (SDN) [20] and network function virtualization (NFV) [21], have significantly enhanced the flexibility and programmability of communication infrastructures supporting ECCC. These technologies facilitate dynamic network resource allocation and traffic engineering, which are essential for optimizing data flows between edge devices and cloud servers [22]. Furthermore, emerging wireless communication paradigms, including millimeter-wave communications, massive MIMO, and non-orthogonal multiple access (NOMA) [23], are revolutionizing the capacity, reliability, and efficiency of edge-cloud connectivity. The integration of network slicing techniques [24] enables the creation of virtualized, isolated network segments with tailored quality-of-service guarantees for diverse edge-cloud applications. Moreover, cross-layer optimization approaches that jointly consider communication, computation, and caching resources [25] are becoming increasingly important for minimizing end-to-end latency and maximizing system throughput. As communication networks evolve toward 6G, the synergistic optimization of networking and computing resources will be instrumental in addressing the growing demands of distributed intelligence across the edge-cloud continuum [26].
@@ I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, yet with varying scopes and technical emphases. Early works like Abbas et al. [27] and Ferrer et al. [28] establish a foundational understanding of edge computing and decentralized cloud architectures respectively, but lack emphasis on AI-driven model optimization. Similarly, Liu et al. [29] provide valuable insights into edge computing systems and tools, while Wang et al. [30] contribute significantly to offloading algorithms, yet neither fully addresses the integration of AI techniques with edge-cloud architectures. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [31] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management but limited attention to model optimization techniques. Domain-specific surveys like Arthurs et al. [32] focus on intelligent transportation systems without broader applicability across diverse sectors. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain underexplored, with Qu et al. [33] providing a dedicated investigation into mobile edge intelligence for LLMs, yet primarily focusing on resource-efficient techniques rather than holistic system integration across the computing continuum. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications but lack comprehensive treatment of model optimization frameworks. Moreover, while Yang et al. [4] extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, their work predominantly emphasizes network dynamics rather than systematically addressing privacy and cross-domain applications. Even the most recent comprehensive works by Yao et al. [11], Gu et al. [34], and Wang et al. [2] exhibit limitations—addressing either model optimization or edge-cloud collaboration but not systematically integrating with comprehensive privacy frameworks and cross-domain applications. As summarized in Table I, these surveys collectively fail to address the full spectrum of ECCC, particularly in integrating model optimization, privacy, security, and diverse applications. Consequently, our survey stands out by providing a holistic analysis that encompasses the entire computing continuum, advanced technical focuses including LLMs, and practical applications across manufacturing, autonomous driving, and healthcare, thereby filling a critical gap in the literature.
@@
Edge Cloud EC Comp. ECC Model Opt. Task Off. Res. Mgmt. Privacy Security LLMs Manuf. Auto. Drive Health
@@ I-C Scope and Contributions
As illustrated in Fig. 1, we provide a systematic and multi-layered examination of ECCC, addressing the critical intersection of distributed intelligence and model optimization. Edge-cloud systems face unprecedented challenges due to heterogeneous infrastructures, resource limitations, and growing computational demands from AI applications, making a comprehensive review of this rapidly evolving field both timely and necessary. By systematically analyzing edge-cloud collaborative computing architectures, resource management strategies, and performance metrics, the survey bridges existing gaps in understanding how AI can enhance decentralized computing environments. Moreover, it investigates critical enabling technologies such as model compression, adaptation, resource-aware optimization, and security mechanisms, which are pivotal for overcoming current limitations. The scope spans from architectural foundations and optimization techniques to practical applications and future research directions, offering researchers and practitioners a holistic perspective of this transformative paradigm. The contributions of this work are:
@@ I-D Research Methodology and Organization
The remainder of this survey is structured to provide a clear and comprehensive overview of AI-driven ECCC. We begin by establishing the fundamentals of edge-cloud computing architecture, including paradigms, distributed intelligence concepts, resources, and metrics in Sec. II. Sec. III examines model optimization techniques including compression, adaptation, resource-aware optimization, and neural architecture search. Subsequently, Sec. IV explores AI-driven resource management covering task offloading, resource allocation, and energy efficiency optimization, while Sec. V addresses privacy and security enhancement strategies. Practical applications appear in Sec. VI through applications spanning autonomous driving, smart healthcare, industrial automation, and smart cities. Sec. VII delivers comprehensive performance analysis through evaluation metrics, benchmarking methods, and system design evaluation, followed by identification of current challenges and future research directions in Sec. VIII, with concluding remarks in Sec. IX.
@@ II Architecture and Fundamentals
Addressing the demands of modern intelligent applications requires architectures that efficiently process vast data volumes with low latency, thus motivating the edge-cloud computing paradigm. Fig. 3 provides a conceptual illustration of such AI-driven edge-cloud collaborative systems, depicting the core infrastructure interplay between cloud servers and edge nodes, alongside essential AI-enabled techniques (e.g., model adaptation, resource-aware optimization) that facilitate distributed intelligence across various application scenarios. Building upon this visual overview, the subsequent subsections delve into the foundational aspects of this architecture. We begin by outlining the core computing paradigm in Sec. II-A, subsequently examining the distributed intelligence framework enabling collaboration in Sec. II-B. Following this, an analysis of crucial resource types and their inherent constraints is presented in Sec. II-C.
@@ II-A Edge-Cloud Computing Paradigm
Edge-cloud computing distributes computation and storage closer to data sources, leveraging a three-tier architecture of cloud, edge, and end devices [38]. The cloud tier provides extensive computational and storage resources for large-scale data processing and model training, while the edge tier, comprising edge servers and gateways, offers localized processing to reduce latency [39]. End devices, such as smartphones and IoT sensors, generate and consume data at the network edge. Effective resource management and orchestration are crucial for efficient resource allocation across these tiers. Distributing intelligence across the network, a key characteristic of this paradigm, can be represented by a function $D(R_{i},L_{j})$, quantifying the proportion of resource $R_{i}$ allocated to layer $L_{j}$ [40]. Edge-cloud systems aim to optimize resource allocation, considering factors like application requirements and network conditions, to maximize a utility function $U(P,C,L)$ encompassing performance $P$, cost $C$, and latency $L$, subject to resource constraints [41]. Consequently, processing data closer to its source reduces latency and bandwidth consumption. Additionally, distributing workloads enhances scalability and fault tolerance, improving the system’s ability to handle increasing data volumes and withstand node failures [42]. Edge federation models further enable collaborative resource sharing, optimizing service delivery across a federated network [43], especially for mission-critical applications [44].

### v4

@@ I-A Background
The increasing prevalence of Internet of things (IoT) devices and the demand for real-time applications have propelled a shift from centralized cloud computing to a distributed edge-cloud paradigm [1, 2, 3]. While cloud computing initially offers scalability and cost-effectiveness, limitations such as latency and bandwidth constraints spur the emergence of edge computing, bringing computation closer to data sources [4, 5, 6]. Consequently, edge-cloud collaborative computing (ECCC) arises, integrating the strengths of both paradigms to facilitate efficient resource utilization and enable complex AI tasks across the edge-cloud continuum [7, 8]. AI integration within ECCC systems addresses the inherent complexity of distributed environments, enabling optimized resource allocation and strategic model deployment for enhanced system performance [9]. In addition, AI empowers intelligent decision-making at the edge, enabling real-time responses. However, realizing distributed intelligence in ECCC systems introduces challenges, including efficient model partitioning and resource-aware model optimization, along with ensuring seamless edge-cloud collaboration and addressing security and privacy concerns [10, 11].
Technical foundations underlying ECCC span multiple critical domains that require comprehensive understanding. Model optimization represents a fundamental challenge in deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency and performance [12]. Model compression techniques including pruning, quantization, and knowledge distillation, alongside adaptive learning approaches such as transfer learning, federated learning, and continual learning enable models to evolve with changing data distributions [6]. Resource management in distributed edge-cloud environments involves complex orchestration of computational, communication, and storage resources across heterogeneous infrastructure, requiring intelligent task offloading strategies, dynamic resource allocation algorithms, and energy-efficient optimization techniques to balance performance, cost, and sustainability [13]. Security and privacy protection mechanisms are paramount in ECCC systems due to the distributed nature of data processing and the involvement of multiple stakeholders, necessitating advanced cryptographic protocols, privacy-preserving machine learning techniques, and secure multi-party computation approaches to protect sensitive information while enabling collaborative intelligence [14].
@@
ECCC has emerged as a crucial paradigm to support the growing demands of real-time, intelligent applications, particularly with the proliferation of artificial intelligence of things (AIoT) devices [15]. However, inherent limitations in edge devices, such as constrained computational power and limited battery life, hinder their ability to independently handle computationally intensive AI workloads [16]. Additionally, the dynamic nature of workloads in these environments requires sophisticated resource allocation and task offloading strategies to maintain performance and prevent bottlenecks [17]. Applications like autonomous driving [18] and industrial automation [19] further necessitate low latency and high reliability, demanding intelligent resource management and optimization [20]. Consequently, AI offers a powerful set of tools, including deep reinforcement learning (DRL) for dynamic resource allocation [21] and multi-objective optimization for task scheduling [22], to address these challenges and improve the efficiency and performance of ECCC systems. For instance, federated learning enables collaborative model training across distributed devices while preserving privacy [16], ultimately contributing to enhanced quality of service (QoS) by optimizing resource utilization and minimizing latency.
@@
Communication network evolution plays a pivotal role in enabling effective edge-cloud collaboration by providing the foundational infrastructure that connects distributed computational resources. Recent advancements in network technologies, particularly software-defined networking (SDN) [23] and network function virtualization (NFV) [24], have significantly enhanced the flexibility and programmability of communication infrastructures supporting ECCC. Advanced technologies facilitate dynamic network resource allocation and traffic engineering, which are essential for optimizing data flows between edge devices and cloud servers [25]. Furthermore, emerging wireless communication paradigms, including millimeter-wave communications, massive MIMO, and non-orthogonal multiple access (NOMA) [26], are revolutionizing the capacity, reliability, and efficiency of edge-cloud connectivity. The integration of network slicing techniques [27] enables the creation of virtualized, isolated network segments with tailored quality-of-service guarantees for diverse edge-cloud applications. Moreover, cross-layer optimization approaches that jointly consider communication, computation, and caching resources [28] are becoming increasingly important for minimizing end-to-end latency and maximizing system throughput. As communication networks evolve toward 6G, the synergistic optimization of networking and computing resources will be instrumental in addressing the growing demands of distributed intelligence across the edge-cloud continuum [29].
The selection and optimization of communication protocols significantly impact the performance, security, and reliability of edge-cloud collaborative systems [30]. Protocol optimization encompasses multiple layers with distinct considerations: (1) Application-layer protocols such as HTTP/HTTPS remain fundamental for web-based edge services, while MQTT and CoAP provide lightweight alternatives specifically designed for resource-constrained IoT devices with intermittent connectivity [31]. MQTT’s publish-subscribe model offers excellent scalability for sensor data aggregation scenarios, whereas CoAP’s REST-based approach provides seamless integration with existing web infrastructures, though MQTT introduces broker dependency that can create single points of failure, while CoAP’s UDP-based nature may require additional mechanisms for reliable delivery in lossy networks [32]; (2) Network-layer protocols including IPv6 and emerging protocols like named data networking (NDN) address the scalability challenges of massive IoT deployments. IPv6 provides abundant address space essential for large-scale edge device connectivity, while NDN’s content-centric approach enables efficient data distribution and caching across edge networks, nevertheless IPv6 adoption remains inconsistent across heterogeneous edge environments, and NDN requires significant infrastructure changes that may limit near-term deployment [33]; and (3) Transport-layer innovations such as QUIC and specialized protocols like BBR congestion control offer improved performance over traditional TCP in high-latency, variable-bandwidth edge-cloud connections [34]. QUIC’s built-in encryption and reduced connection establishment overhead particularly benefit mobile edge scenarios, while BBR’s model-based congestion control adapts better to the diverse network conditions encountered in edge deployments. The choice of appropriate protocols depends on specific application requirements, device capabilities, and network characteristics, with optimal edge-cloud systems often employing protocol stacks that dynamically adapt to changing conditions and requirements [35].
@@ I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, each with distinct scopes and technical emphases. Early works like Abbas et al. [36] and Ferrer et al. [37] establish a foundational understanding of edge computing and decentralized cloud architectures respectively, with their primary focus on architectural foundations rather than AI-driven model optimization. Similarly, Liu et al. [38] provide valuable insights into edge computing systems and tools, while Wang et al. [39] contribute significantly to offloading algorithms, with their emphasis primarily on different aspects of edge-cloud integration rather than comprehensive AI technique integration. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [40] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management, with their primary emphasis on system-level considerations rather than model optimization techniques. Domain-specific surveys like Arthurs et al. [41] focus on intelligent transportation systems with specialized depth in this domain. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain an active area of exploration, with Qu et al. [42] providing a dedicated investigation into mobile edge intelligence for LLMs, focusing primarily on resource-efficient techniques within their specialized scope. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications with their emphasis on domain-specific considerations.
The most relevant contemporary works include Yang et al. [4], who extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, with their work emphasizing network dynamics as their primary contribution. Similarly, the comprehensive works by Yao et al. [11], Gu et al. [43], and Wang et al. [2] each make valuable contributions with distinct emphases. Specifically, [11] provides excellent coverage of edge-cloud collaboration and AI techniques with particular strength in these core areas. [43] delivers strong analysis of AI-enhanced cloud-edge-terminal networks with good application coverage, emphasizing network architecture aspects. [2] offers comprehensive model optimization coverage across the end-edge-cloud spectrum with particular depth in optimization techniques.
Two surveys are particularly relevant to our work and warrant detailed comparison. Yang et al. [4] provides a comprehensive survey of reinforcement learning approaches for mobile edge computing with excellent coverage of resource management and autonomous driving applications. Their work has a specific focus that differs from our survey in several key aspects: (1) Specialized computing paradigm emphasis, while they focus extensively on edge computing and resource management, our work provides comprehensive coverage of cloud integration and edge-cloud collaboration frameworks; (2) Domain-specific technical focus, their emphasis on reinforcement learning, while thorough, represents a specialized approach compared to our comprehensive treatment of model optimization techniques such as compression, adaptation, and neural architecture search that are central to modern ECCC systems; (3) Different scope for privacy and security, their work emphasizes other technical aspects, while our survey provides systematic attention to privacy-preserving mechanisms and security enhancement for real-world ECCC deployments; (4) Specialized application coverage, while strong in autonomous driving, our work provides comprehensive coverage of manufacturing and healthcare applications that increasingly rely on edge-cloud collaboration. Qu et al. [42] presents a focused investigation into mobile edge intelligence for large language models with strong coverage of LLMs, privacy considerations, and healthcare applications. Their work complements our survey with different emphases: (1) Specialized technical scope, while excellent in LLM-specific optimizations, our survey provides systematic coverage of general model optimization techniques applicable across diverse AI models and architectures; (2) Mobile edge intelligence focus, their focus on mobile edge intelligence represents a specialized area, while our work emphasizes comprehensive edge-cloud collaboration frameworks and distributed intelligence paradigms; (3) Specialized application domain, while strong in healthcare applications, our survey provides detailed analysis of manufacturing and industrial IoT scenarios that represent major application domains for ECCC; (4) Domain-specific security analysis, their privacy coverage addresses specific requirements, while our work provides comprehensive treatment of security mechanisms and adversarial considerations essential for robust ECCC systems.
In contrast to these works, our survey uniquely provides: (1) Comprehensive computing continuum coverage spanning edge, cloud, edge-cloud computing, and collaborative frameworks; (2) Holistic technical integration systematically covering model optimization (compression, adaptation, NAS), resource management, privacy, security, and emerging LLM paradigms; (3) Diverse application analysis with detailed coverage across manufacturing, autonomous driving, and healthcare domains; (4) Unified framework approach that synthesizes distributed intelligence principles with practical implementation considerations. As summarized in Table I, these surveys collectively address different aspects of ECCC, with each making valuable contributions to specific areas of the field. Our survey builds upon these foundational works by providing integrated coverage of model optimization, privacy, security, and diverse applications within a unified framework. Most importantly, existing works have established strong foundations in specific technical aspects (e.g., model optimization or resource management) or particular application domains (e.g., transportation or healthcare), while our work provides a holistic synthesis that spans the entire computing continuum. Consequently, our survey contributes by providing comprehensive analysis that encompasses advanced technical focuses including LLMs, and practical applications across manufacturing, autonomous driving, and healthcare, thereby complementing and extending the existing literature.
@@
Edge Cloud ECCo ECC Model Opt. Task Off. Res. Mgmt. Privacy Security LLMs Manuf. Auto. Drive Health
@@ I-C Scope and Contributions
As illustrated in Fig. 1, we provide a systematic and multi-layered examination of ECCC, addressing the critical intersection of distributed intelligence and model optimization. Edge-cloud systems face unprecedented challenges due to heterogeneous infrastructures, resource limitations, and growing computational demands from AI applications, making a comprehensive review of this rapidly evolving field both timely and necessary.
Our survey establishes a comprehensive taxonomy that organizes ECCC into five interconnected layers, each addressing specific aspects of distributed intelligence deployment. Foundations layer encompasses architectural paradigms, distributed intelligence frameworks, and fundamental resource constraints that govern system design. Technical implementation layer focuses on core enabling technologies including model optimization techniques (compression, adaptation, neural architecture search) and AI-driven resource management strategies (task offloading, resource allocation, energy optimization). Protection framework layer addresses critical security and privacy concerns through advanced cryptographic protocols, privacy-preserving learning, and secure computation techniques. Practical deployment layer examines real-world applications across diverse domains and comprehensive performance analysis methodologies. Finally, the Research vision layer identifies current challenges and future research directions that will shape the evolution of ECCC systems. Our taxonomic framework provides researchers and practitioners with a structured understanding of how different components interact within the broader ECCC ecosystem, enabling systematic development and optimization of distributed intelligent systems.
By systematically analyzing ECCC architectures, resource management strategies, and performance metrics, the survey bridges existing gaps in understanding how AI can enhance decentralized computing environments. Moreover, it investigates critical enabling technologies such as model compression, adaptation, resource-aware optimization, and security mechanisms, which are pivotal for overcoming current limitations. The scope spans from architectural foundations and optimization techniques to practical applications and future research directions, offering researchers and practitioners a holistic perspective of this transformative paradigm.
The three primary technical categories of ECCC each address distinct challenges and exhibit specific applicability constraints. Model optimization techniques (Sec. III) primarily target the fundamental challenge of deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency. Model optimization approaches are particularly suitable for scenarios involving deep learning model deployment, privacy-sensitive distributed training, and real-time inference applications. However, their effectiveness is constrained by the inherent trade-offs between model accuracy and compression ratios, the heterogeneity of edge device capabilities, and the complexity of managing distributed learning processes. Model compression methods excel in reducing computational overhead but may compromise accuracy, while federated learning approaches provide privacy preservation at the cost of increased communication complexity and convergence challenges with non-IID data distributions. AI-driven resource management strategies (Sec. IV) focus on optimizing resource utilization and task orchestration across the heterogeneous edge-cloud continuum. Resource management approaches are most effective in scenarios with dynamic workloads, multi-tenant environments, and applications requiring intelligent load balancing across diverse computational resources. The primary limitations include the computational complexity of multi-objective optimization problems, the difficulty in predicting system behavior under varying conditions, and scalability challenges in large-scale deployments. While task offloading mechanisms can significantly reduce latency and energy consumption, they require sophisticated algorithms to handle dependency relationships and may introduce communication overhead that negates performance benefits. Privacy and security enhancement frameworks (Sec. V) address the critical requirements for protecting sensitive data and ensuring system trustworthiness in distributed environments. Security techniques are essential for healthcare, financial, and government applications where regulatory compliance and data protection are paramount. Nevertheless, privacy-preserving mechanisms often introduce computational overhead and complexity in key management, while security measures must continuously evolve to address emerging threats in heterogeneous edge environments.
@@ I-D Research Methodology and Organization
The remainder of this survey provides a comprehensive overview of AI-driven ECCC. We begin with edge-cloud computing fundamentals including architecture and distributed intelligence in Sec. II. Sec. III examines model optimization techniques including compression and neural architecture search. Sec. IV explores AI-driven resource management covering task offloading and energy efficiency, while Sec. V addresses privacy and security enhancement. Sec. VI presents practical applications spanning autonomous driving, healthcare, and smart cities. Sec. VII delivers performance analysis through benchmarking and evaluation, followed by challenges and future directions in Sec. VIII, with conclusions in Sec. IX.
@@ II Architecture and Fundamentals
Modern intelligent applications demand architectures capable of processing vast data volumes with minimal latency, driving innovation toward distributed computing paradigms that bridge resource-rich cloud infrastructure with proximity-based edge processing. Fig. 3 provides a conceptual illustration of AI-driven edge-cloud collaborative systems, depicting the core infrastructure interplay between cloud servers and edge nodes, alongside essential AI-enabled techniques (e.g., model adaptation, resource-aware optimization) that facilitate distributed intelligence across various application scenarios. Building upon this visual overview, the subsequent subsections delve into the foundational aspects of this architecture. We begin by outlining the core computing paradigm in Sec. II-A, subsequently examining the distributed intelligence framework enabling collaboration in Sec. II-B. Following this, an analysis of crucial resource types and their inherent constraints is presented in Sec. II-C.
@@ II-A Edge-Cloud Computing Paradigm
Distributed computing architectures optimize performance by positioning computational and storage resources closer to data sources, implementing a three-tier hierarchy encompassing cloud infrastructure, edge nodes, and end devices [47]. The cloud tier provides extensive computational and storage resources for large-scale data processing and model training, while the edge tier, comprising edge servers and gateways, offers localized processing to reduce latency [48]. End devices, such as smartphones and IoT sensors, generate and consume data at the network edge. Effective resource management and orchestration are crucial for efficient resource allocation across these tiers. Distributing intelligence across the network, a key characteristic of this paradigm, can be represented by a function $D(R_{i},L_{j})$, quantifying the proportion of resource $R_{i}$ allocated to layer $L_{j}$ [49]. Edge-cloud systems aim to optimize resource allocation, considering factors like application requirements and network conditions, to maximize a utility function $U(P,C,L)$ encompassing performance $P$, cost $C$, and latency $L$, subject to resource constraints [50]. Consequently, processing data closer to its source reduces latency and bandwidth consumption. Additionally, distributing workloads enhances scalability and fault tolerance, improving the system’s ability to handle increasing data volumes and withstand node failures [51]. Edge federation models further enable collaborative resource sharing, optimizing service delivery across a federated network [52], especially for mission-critical applications [53].